OpenStudy (anonymous):

I really need help FAST!! Which set of fractions is ordered from least to greatest? A two over five, one over four, three over eight B three over eight, two over five, one over four C one over four, three over eight, two over five D three over eight, one over four, two over five

OpenStudy (texaschic101):

2/5 , 1/4 , 3/8 --- common denominator is 40 2/5 = 16/40 1/4 = 10/40 3/8 = 15/40 now put them from least to greatest

OpenStudy (anonymous):

Find A common denominator...

OpenStudy (anonymous):

WOW, ok so how did u get that. Ive been out of school for 15 years, and all if this stuff is like NEW to me O_o Ima bout to pull my hair out LOL

OpenStudy (anonymous):

Ok heres another one: IM SO THANKFUL YOU GUYS!!, if u only new. My 25 year old daughter told me to come on here and she said i was bound to get help from someone! Which set of fractions is ordered from least to greatest? A seven over eight, five over eleven, two over three B five over eleven, seven over eight, two over three C two over three, five over eleven, seven over eight D five over eleven, two over three, seven over eight

OpenStudy (texaschic101):

you can do this 2 ways...you can either turn them into fractions with a common denominator, then compare them. Or, you can turn them all into decimals and compare them....whichever is easier for you to do

OpenStudy (texaschic101):

7/8....grab the calculator = 0.875 5/11 = 0.4545 2/3 = 0.666 now put them in the order you need them in

OpenStudy (anonymous):

ok for the second on lets just sa the fraction is 2/3 so would is say 3 divided by 2

OpenStudy (texaschic101):

punch in 2 , then division sign, then 3

OpenStudy (anonymous):

oh ok! Im exceited Im learning. Im chatting with my daughter on IM while shes at work, she is shaking her head at me LOL,

OpenStudy (texaschic101):

lol....you will soon get it

OpenStudy (anonymous):

ok im gonna try the next one by myself. brb! heres the question ... Which set of fractions is ordered from greatest to least? A two over three, three over eight, five over twelve B five over twelve, three over eight, two over three C two over three, five over twelve, three over eight D three over eight, two over three, five over twelve

OpenStudy (texaschic101):

its never too late to learn

OpenStudy (skullpatrol):

Binging excited about learning deserves a medal +1

OpenStudy (anonymous):

is it C!

OpenStudy (texaschic101):

2/3 = 0.666 3/8 = 0.375 5/12 = 0.416 greatest to least : 2/3, 5/12, 3/8...yep, you got it, C is the answer

OpenStudy (anonymous):

(dancing) THANKS!!

OpenStudy (texaschic101):

no problem....I will do the happy dance with you :) whoop - whoop

OpenStudy (anonymous):

KMSL!!! You may not believe me but im taking notes!! LOL Ok now they've moe me on to something else...heres the question...Ive spent so much time on these questions. I have 19 minutes left and 17 questions still :-( Choose the fraction that goes in the blank. six over ten >_______> one over two A one over two B one over four C two over three D three over four

OpenStudy (texaschic101):

do you know what these signs mean < and > ??

OpenStudy (anonymous):

im just guessing is it C?

OpenStudy (anonymous):

yep Greater than and less than

OpenStudy (anonymous):

i may have gotten that one wrong bcuz i just picked one. Im running out of time and I really have to have notpicked anything :-(

OpenStudy (anonymous):

OMG 10 minutes and 11 questions left.... I BOMBED!

OpenStudy (texaschic101):

6/10 = 0.6 1/2 = 0.5 so basically, you need a fraction that is LESS THEN 6/10 (or 0.6) and GREATER THEN 1/2 (or 0.5) 2/3 = 0.66 3/4 = 0.75 1/2 = 0.50 1/4 = 0.25 hmm..not sure ....none seem to match

OpenStudy (anonymous):

Im not sure, ill have to go back at the end and look. how would i solve this one Which decimal represents the fraction of plants in this group that have flowers? Image of 5 plants. 4 of the plants have flowers. A 0.8 B 0.9 C 1.5 D 4.5

OpenStudy (texaschic101):

part / whole = 4/5 = 0.8

OpenStudy (anonymous):

GEESH!! Ok ur gonna have to tell me how u got that, when i finish LOL I never wouldve got that

OpenStudy (anonymous):

Which of the following integers is least? A -5 + (-2) B -5 ÷ (-2) C -5 × (-2) D -5

OpenStudy (anonymous):

D?

OpenStudy (texaschic101):

-5 + (-2) = -5 - 2 = -7 -5 / -2 = -5 * -1/2 = 5/2 = 2.5 -5 * -2 = 10 -5 your least interget would be A... -7

OpenStudy (texaschic101):

typo..integer

OpenStudy (anonymous):

i got that one wrong then

OpenStudy (texaschic101):

you have to watch out for negatives....they are kinda the opposite of positives in the fact that the bigger negative number is actually smaller. -7 is smaller then -5

OpenStudy (anonymous):

Which is true? A 4.02 > 4.1 B 4.181 < 4.118 C 4.5 > 4.49 D 4.665 > 4.67

OpenStudy (texaschic101):

when adding.subtracting fractions, you have to have a common denominator. when dividing fractions, you " flip " what you are dividing by, then multiply. when multiplying fractions, you just multiply straight across

OpenStudy (texaschic101):

C. 4.5 > 4.49

OpenStudy (anonymous):

ok I ran out of time :-( So how did u get the last answer?

OpenStudy (texaschic101):

you basically follow the numbers from the right to the left 4.02 and 4.1 ....they both start with a 4....go to the next number....0 and 1....0 is smaller, so 4.02 < 4.1 4.5 and 4.49....they both start with a 4...go to next number...5 and 4....5 is greater, so 4.5 > 4.49 you kinda understanding ? Don't forget to pay attention to the number of the left of the decimal....for example : 23.3 and 2.33....the number to the right of the decimal is a whole number, so 23.3 > 2.33 because 23 is greater then 2

OpenStudy (texaschic101):

oops...follow the numbers from the left to the right

OpenStudy (texaschic101):

typo...the number to the left of the decimal is a whole number and the numbers to the right of the decimals are less then 1

OpenStudy (anonymous):

lol, its ok I knew what u meant

OpenStudy (texaschic101):

I got to get my directions, left and right, straightened out...lol

OpenStudy (anonymous):

so on the left its like ones, tens, hundreds...and so forth?

OpenStudy (texaschic101):

yes...to the right of the decimal, it is ones, tens, hundreds...etc to the left of the decimal. it starts at tenths, hundredths, thousandths

OpenStudy (anonymous):

Ok "wayment" and my daughter would say LOL.... lets say the number is 1234.5678 so the 4 is ones and the 5 is tenths?

OpenStudy (texaschic101):

yes...the 4 is ones,the 2 is tens, the 5 is tenths, the 6 is hundrdths this might help.. http://www.mathsisfun.com/decimals.html

OpenStudy (texaschic101):

oops..the 3 is tens

OpenStudy (texaschic101):

I would stay and help you more, but I have to go now. I wish you the best of luck

OpenStudy (anonymous):

I totally appreciate the time that you've taken out with me! YOU ROCK!!

OpenStudy (texaschic101):

lol...I don't know about that...but thank you :)

OpenStudy (texaschic101):

your daughter is gonna be proud of you :)

OpenStudy (anonymous):

Yep, she is! I cant teach her something now! LOL...Have a wonderful day Texaschic!